OpenStudy (anonymous):

ok, need to write 81^-3/4 in radical notation. i came up with 4 OVER 3(sqroot)81 and then I have to simplify that to a fraction or integer @satellite73 How'd I do?

OpenStudy (anonymous):

you have an exponent with three feature a) a denominator of 4 b) a numerator of 3 c) a minus sign they mean in turn a) the fourth root b) the cube c) the reciprocal

OpenStudy (anonymous):

so \[81^{-\frac{3}{4}}\] can be written as \[\frac{1}{\sqrt[4]{81^3}}\]

OpenStudy (anonymous):

and evaluated via \[\sqrt[4]{81}=3\] \[3^3=27\] \[\frac{1}{27}\] as your final answer

OpenStudy (anonymous):

I forgot the 1! lol i had it half right... haha. a little mixed up ..but i'm understanding it. thak you :)

OpenStudy (anonymous):

yw